OpenStudy (anonymous):

Given the triangle below, which of the following is a correct statement? Right triangle ABC with AB measuring 4, AC measuring 5, and BC measuring 6. Answer Choices: csc right triangle C right triangle equals fraction 3 over 2 sec right triangle C right triangle equals fraction 5 over 6 csc right triangle B right triangle equals fraction 3 over 2 sec right triangle B right triangle equals fraction 6 over 5

OpenStudy (anonymous):

@ribhu

OpenStudy (ribhu):

diagram

OpenStudy (ribhu):

i don't think this will be a right triangle

OpenStudy (mathstudent55):

Correct. 4, 5, 6 is not a Pythagorean triple, so there is no right triangle.

OpenStudy (ribhu):

this is never a right triangle @sarahnichols3

OpenStudy (anonymous):

Wrong question!! So sorry give me a second!

OpenStudy (anonymous):

@ribhu & @mathstudent55

OpenStudy (mathstudent55):

No statement given is correct, including the statement of the problem that states this is a right triangle.

OpenStudy (mathstudent55):

Ok, post the correct one.

OpenStudy (anonymous):

Actually it was the correct question, here is the print screen of the problem.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@ribhu

OpenStudy (ribhu):

its a big attachment

OpenStudy (anonymous):

is that a problem? @ribhu

OpenStudy (ribhu):

no just internet is a bit slow in here

OpenStudy (anonymous):

My bad!

OpenStudy (ribhu):

wait its downloading

OpenStudy (anonymous):

Okay!

OpenStudy (mathstudent55):

I still say the problem is wrong. A right triangle cannot have sides measuring 4, 5, and 6 units.

OpenStudy (anonymous):

So what do I choose???

OpenStudy (ribhu):

its not possible but if answer is to be marked it would be A

OpenStudy (anonymous):

Okay thank you !

OpenStudy (ribhu):

cosec C =6/4 =3/2

OpenStudy (anonymous):

Thank you!